Problem Solving: How to solve:
A factory employs skilled laborers at $25 per hour and unskilled at $12 per hour. The hourly cost of the factory's 24 workers is $405. How many of each type of laborer works at the factory?
s-skilled
u-unskilled

Answer by checkley71(8403)

25X+12(24-X)=405
25X+288-12X=405
13X=405-288
13X=117
X=117/13
X=9 SKILLED WORKERS.
24-9=15 UNSKILLED WORKERS.
PROOF
25*9+12*15=405
225+180=405
405=405